# A Bayesian hierarchical model for methane emission source apportionment

### v1

@@ Abstract
Keywords: climate change, oil and gas emissions, Bayesian inverse modeling, Gibbs sampling, spike-and-slab prior
@@ 1 Introduction
By 2050, climate change is projected to cause an additional 14.5 million deaths and \$12.5 trillion in economic losses worldwide due to an increase in the frequency and severity of extreme weather events (World Economic Forum, 2024). Rising concentrations of carbon dioxide and methane in the atmosphere are the primary drivers of climate change, making it crucial to reduce emissions of these gasses. Methane, in particular, provides an important opportunity to mitigate the near-term impacts of climate change given its relatively short lifetime in the atmosphere and high heat trapping potential compared to carbon dioxide (Szopa et al., 2021).
The oil and gas industry accounts for 22% of global anthropogenic methane emissions (O’Rourke et al., 2021; Crippa et al., 2021) and is a viable opportunity for emissions reductions given that: 1) emissions from equipment malfunctions can often be mitigated quickly once they are identified, and 2) technological improvements and facility design innovations can eliminate emissions from normally operating equipment such as pneumatic controllers (Nisbet et al., 2020). Knowing which pieces of equipment are emitting and their emission rates at any given time is necessary to both prioritize mitigation efforts and verify that long term emission reduction strategies are working. However, estimating emission source and rate based on ambient methane measurements is a challenging inverse problem due to the turbulent nature of near-field atmospheric transport, the degree of variability in emission rates, and the fact that only a subset of oil and gas equipment are emitting at a given time.
In this article, we analyze in situ sensor data from an experimental oil and gas site where methane is released at known rates in a way that emulates real-world emission characteristics. This site is called the Methane Emissions Technology Evaluation Center (METEC). Sensor data collected at METEC come as close as possible to data from a real oil and gas site while still knowing ground truth emission source and rate. We develop a Bayesian hierarchical model for performing methane emission source apportionment at METEC, and by extension, on real oil and gas sites. We call this model the multisource detection, localization, and quantification (MDLQ) model. The MDLQ model combines an atmospheric transport model with a multivariate time series model for the sensor data. This is the first study, to our knowledge, that pairs a novel inversion method with a controlled release experiment that is sufficiently realistic to generalize to operating oil and gas sites.
@@ 1.1 Methane releases at METEC
The METEC facility, shown in Figure 1, is instrumented with real oil and gas equipment, such as wellheads and separators, but does not actually produce oil or gas. Instead, known quantities of methane are released from typical leak points on the oil and gas equipment. The controlled releases at METEC emulate the following characteristics of methane emissions on active oil and gas sites: emissions can last from minutes to days with rates that vary over time, any subset of the equipment can emit simultaneously, and there can be periods of no emissions (Cusworth et al., 2021; Wang et al., 2022; Daniels et al., 2023).
@@
The METEC experiment consists of 337 releases conducted from February 5 to April 30, 2024. Releases are separated by periods of no emissions, and each release consists of emissions from a subset of the five equipment groups shown in Figure 1. These experiments provide a significantly larger and more realistic sample of methane emissions than other site-level inversion studies in the literature. For example, Cartwright et al. (2019) analyze two releases, Newman et al. (2024) analyze three releases, and Kumar et al. (2022) analyze 26 releases. The ability to apportion methane emissions using a small number of sensors is influenced by site layout, wind, and other meteorological conditions, making it important to assess the performance of a statistical inversion over a wide range of conditions.
Furthermore, the releases in this study better emulate the emission characteristics of active oil and gas sites than similar studies in the literature. For example, Weidmann et al. (2022) evaluate their method on methane releases with known start and end times, and Hirst et al. (2020) perform releases in flat empty locations with simple atmospheric transport. As such, the modeling results in this study are more likely to generalize to real oil and gas sites of similar size and complexity to METEC and provide a baseline level of apportionment accuracy on these sites.
@@ 1.2 The continuous monitoring inverse problem
The MDLQ model uses methane concentration measurements from a network of in situ sensors placed directly on oil and gas sites to apportion emissions. We refer to this type of sensor system as a continuous monitoring system (CMS). The locations of the CMS sensors in this study are shown as pins in Figure 1. Each sensor measures ambient methane concentrations once per second, which are averaged in situ every minute. Let $\boldsymbol{y}\equiv\{y_{1},...,y_{n}\}$ denote the minute-frequency CMS concentration measurements, where $n$ is determined by the amount of time over which the inversion is to be taken (discussed further in Section 3.3). As with most remote sensing technologies, an atmospheric inversion is required to translate $\boldsymbol{y}$ into estimates of emission source and rate.
@@
where emission rates for the $p$ potential sources on the site, $\boldsymbol{\beta}\equiv\{\beta_{1},...,\beta_{p}\}$, are assumed to be constant within the temporal extent of the inversion. The design matrix is often constructed as $X\equiv[\boldsymbol{x}^{(1)}\cdot\cdot\cdot\boldsymbol{x}^{(p)}]\in\mathbb{R}^{n\times p}$, where each $\boldsymbol{x}^{(i)}\in\mathbb{R}^{n}$ for $i=\{1,...,p\}$ contain the concentration predictions from the transport model at all sensor locations assuming that source $i$ is emitting at a unitary rate. In many studies, the errors $\boldsymbol{\epsilon}\equiv\{\epsilon_{1},...,\epsilon_{n}\}$ are assumed normal, independent, and identically distributed (Ars et al., 2017; Cartwright et al., 2019; Hirst et al., 2020; Newman et al., 2024), which is another limitation given that $\boldsymbol{y}$ are time series data. In this study, we are primarily concerned with estimates for $\boldsymbol{\beta}$, as these tell us the subset of sources that are emitting and their rate.
Our approach, implemented in the MDLQ model, assumes the same linear model as Equation 1 at the data level, but further models $\boldsymbol{\beta}$ as a mixture distribution with a point mass at 0 kg/hr. This process model allows for rate estimates to be identically zero, a necessary property given that potential sources are often not emitting. We use a Dirac delta function to implement the point-mass, providing an extension to the methodology in Weidmann et al. (2022), who used a narrow Gaussian to approximate the point mass. This extension eliminates the need to tune the relative widths of the components of the mixture distribution. Finally, following Ganesan et al. (2014), we allow $\boldsymbol{\epsilon}$ to be autocorrelated since the elements of $\boldsymbol{y}$ are observed over time at 1-minute intervals. We extend the methods in Ganesan et al. (2014) by parameterizing the covariance matrix such that the determinant and inverse have closed forms, allowing for fast computation.
@@ 2.1 METEC controlled release and measurement data
Methane concentration data for the METEC experiment come from a network of 10 CMS sensors installed around the facility at a height of 2 m. Eight of the sensors are also equipped with an anemometer to measures wind speed and direction in the horizontal plane. Wind data are necessary to forward simulate the transport of methane from the emission sources to the sensor locations (discussed further in Section 2.2). The properties (e.g., accuracy and precision) of the CMS sensors in this study are comparable to the sensors being used operationally to mitigate emissions. Full sensor specifications are provided in Section S1 of the Supporting Information (SI) file.
Unlike many studies in the literature, the controlled releases in this experiment were performed around-the-clock to evaluate the model under both daytime and nighttime meteorology. Each release consists of emissions from a subset of the five potential emission sources shown in Figure 1. Emission rates were constant within a given release, but could be different between the emitting equipment. Across the 337 releases, equipment-level release rates ranged from 0.08 to 7.2 kg/hr, emission durations ranged from 1.1 minutes to 8.0 hours, and release heights ranged from 0.12 m to 5.5 m.
@@
As an example, Figure 2 shows the true emission state during a 24-hour period and the corresponding concentration and wind data from the CMS sensors. In Figure 2(b), we plot the minute-by-minute maximum concentration observation across the 10 CMS sensors as a condensed summary of the data. Figure S4 in the SI shows the methane concentration measurements from all 10 CMS sensors separately. Notice that the concentration enhancements align with the timing of the controlled releases, but the amplitude of the enhancements do not follow the magnitude of the releases. This is because the sensors are not located at the release points, and meteorological effects such as wind speed affect the transport of methane and the resulting concentrations at the sensor locations. Modeling this transport is critical for an accurate inversion of emission source and rate.
@@ 2.2 Atmospheric transport model
Consider a coordinate system in which emission source $i$ is located at $(0,0,H_{i})$. The puffs of methane from this source will all originate at this point. New puffs are created at a given frequency, $\delta t$, by default once per second, and the coordinates are rotated for each puff so that the positive $x$-direction is pointed in the downwind direction at the time of puff creation. In this coordinate system, the predicted methane concentration at sensor location $(x,y,z)$ from a given puff, index by $j$, that originated at $(0,0,H_{i})$ and has been in existence for time $t$ is modeled as
$c_{j}^{(i)}(x,y,z,t)=\frac{Q}{(2\pi)^{3/2}\sigma_{y}^{2}\sigma_{z}}\exp{\left(-\frac{(x-ut)^{2}+y^{2}}{2\sigma_{y}^{2}}\right)}\left[\exp{\left(-\frac{(z-H_{i})^{2}}{2\sigma_{z}^{2}}\right)}+\exp{\left(-\frac{(z+H_{i})^{2}}{2\sigma_{z}^{2}}\right)}\right],$
where $Q$ is the mass of methane contained in puff $j$, $\sigma_{y}$ and $\sigma_{z}$ are the dispersion parameters in the $y$ and $z$ directions, respectively, and $u$ is the wind speed at the time of puff creation. The value of $Q$ is a function of the specified emission rate, $q$, and the puff creation frequency, $\delta t$. The GPM requires time, $t$, to be discretized at a fixed multiple of the puff creation frequency, $\delta t$.
At each subsequent time step, a new puff is created at $(0,0,H_{i})$ and puffs already in existence are transported based on the wind speed and direction from their respective creation times. The GPM used in this study assumes zero advection in the vertical direction, as wind data in this dimension are hard to obtain in practice, but the puffs do grow more diffuse in all three physical dimension as they move downwind. The total methane concentration at sensor location $(x,y,z)$ and time $t$ from source $i$ is then
$c^{(i)}(x,y,z,t)=\sum_{j=1}^{P}c_{j}^{(i)}(x,y,z,t),$
where $P$ is the number of puffs in existence at time $t$.
@@ 2.3 Background estimation and removal
This method for estimating the methane background imposes no assumptions about its spatial or temporal homogeneity nor does it require fitting a spatiotemporal model. Furthermore, we have found that estimating the background using a fixed value, such as ${}^{\text{th}}$ percentile from Cartwright et al. (2019), often underestimates the true background by a few parts per million. This can lead to large errors in the inferred emission rates when the amplitude of the concentration enhancements is on a similar scale as the errors in the background estimate.
@@ 3.1 Bayesian hierarchical model
There are three layers to the MDLQ model, which are summarized in Figure 4. The data layer (highest layer) contains a model for the CMS concentration observations. The process layer (middle layer) contains a model for the latent methane emission process that we do not directly observe. Finally, the parameter layer (lowest layer) contains priors for the model parameters that are not of direct interest for methane emission source apportionment but still contain important uncertainty information.
@@
3.1.1 Data layer
Our model for the CMS concentration data takes the form of a multiple linear regression,
$\boldsymbol{y}=X\boldsymbol{\beta}+\boldsymbol{\epsilon},\vspace{-5mm}$
where $\boldsymbol{y}\equiv\{y_{1},...,y_{n}\}$ are the background-corrected CMS concentration observations at a 1-minute frequency. We “stack” the concentration observations from all 10 CMS sensors in $\boldsymbol{y}$, as we want to leverage the information provided by each sensor jointly when performing the inversion, rather than performing a separate inversion for each sensor. The length of $\boldsymbol{y}$ will reflect the choice of inversion length (discussed in Section 3.3), as $n$ is equal to the number of 1-minute time steps used in the inversion multiplied by the number of sensors. We define the design matrix as $X\equiv[\boldsymbol{x}^{(1)}\cdot\cdot\cdot\boldsymbol{x}^{(p)}]\in\mathbb{R}^{n\times p}$, where each $\boldsymbol{x}^{(i)}\in\mathbb{R}^{n}$ for $i=\{1,...,p\}$ denotes the concentration predictions from the GPM at all 10 CMS sensor locations assuming that source $i$ is emitting at a unitary emission rate. We again “stack” the predicted concentrations at all sensor locations in $\boldsymbol{x}^{(i)}$ to match the form of $\boldsymbol{y}$. Using the notation from Section 2.2, each $\boldsymbol{x}^{(i)}=[c^{(i)}(x_{1},y_{1},z_{1},t\in T),...,c^{(i)}(x_{10},y_{10},z_{10},t\in T)]$, where $x_{k}$, $y_{k}$, and $z_{k}$ are the spatial location of the $k^{\text{th}}$ sensor, $T$ is the temporal domain of the inversion, and $c(\cdot)$ are the concentration predictions assuming that source $i$ is emitting. Finally, $\boldsymbol{\beta}\equiv\{\beta_{1},...,\beta_{p}\}$ are the emission rates for the $p$ sources on the site, which are assumed to be constant within the temporal extent of the inversion.
We assume that the errors $\boldsymbol{\epsilon}\equiv\{\epsilon_{1},...,\epsilon_{n}\}$ follow a multivariate normal, such that $\boldsymbol{\epsilon}\sim N(0,\sigma^{2}R)$, where $R\in\mathbb{R}^{n\times n}$ contains the correlation structure of an AR(1) process and is solely a function of a correlation parameter $r\in[0,1]$. The exact form of $R$ is given in Section S10 of the SI. Modeling autocorrelation in $\boldsymbol{\epsilon}$ extends previous work and results in better coverage of the credible intervals (see Section S8 in the SI for details). The order of the autocorrelation model was selected through a visual analysis of partial autocorrelation functions; we find that an AR(1) often captures most of the autocorrelation in $\boldsymbol{\epsilon}$. Additionally, using an AR(1) results in a covariance matrix with closed forms for both the inverse and determinant (Bodin et al., 2012), which allow for efficient sampling.
3.1.2 Process layer
@@
The methane emission rate for each source takes the form of a mixture model with a point mass at 0 kg/hr, given by
$\beta_{i}$$\sim\begin{cases}0&z_{i}=0\\ \text{Exp}(\tau_{i}^{2}\sigma^{2})&z_{i}=1,\end{cases}$
@@
where the indicator $z_{i}\in\{0,1\}$ is modeled as $z_{i}\sim\text{Bernoulli}(\theta_{i})$. This model is an example of a sparsity enforcing spike-and-slab prior (Mitchell & Beauchamp, 1988). Sparsity in the emission rates is an intentional property of the MDLQ model, as potential emission sources on oil and gas sites are often not emitting. Previous work has found that emission rates over time follow a right skewed distribution (Daniels et al., 2023), so we use an exponential for the non-zero component of the mixture parameterized with a scale parameter given by $\tau_{i}^{2}\sigma^{2}$. This scale parameter allows the exponential to scale with the magnitude of $\boldsymbol{y}$, making it easier to interpret MDLQ parameter estimates in other applications where $\boldsymbol{y}$ might have a very different magnitude. Furthermore, the exponential component forces $\boldsymbol{\beta}$ to have non-negative support, as there are no methane sinks on oil and gas sites.
@@
This variant of the spike-and-slab prior uses a discrete point-mass as the spike rather than a continuous distribution, such as an exponential with scale parameter much smaller than $\tau_{i}^{2}\sigma^{2}$. This allows for $\beta_{i}$ to be sampled as identically zero, rather than a very small value. Furthermore, it eliminates the need to prescribe or infer the scale parameter of the narrow distribution approximating the discrete point-mass, which can be challenging to do in such a way that the MCMC sampler is efficient and mixes well.
3.1.3 Parameter layer

### v2

@@ Abstract
spike-and-slab prior,
keywords:
[…]
and
@@ 1 Introduction
By 2050, climate change is projected to cause an additional 14.5 million deaths and \$12.5 trillion in economic losses worldwide due to an increase in the frequency and severity of extreme weather events (World Economic Forum, 2024). Rising concentrations of carbon dioxide and methane in the atmosphere are the primary drivers of climate change, making it crucial to reduce emissions of these gasses (IPCC, 2021). Methane, in particular, provides an important opportunity to mitigate the near-term impacts of climate change given its relatively short lifetime in the atmosphere and high heat trapping potential compared to carbon dioxide (Szopa et al., 2021).
The oil and gas industry accounts for 21% of global anthropogenic methane emissions (International Energy Agency, 2025) and is a viable opportunity for emissions reductions given that: 1) emissions from equipment malfunctions can often be mitigated quickly once they are identified, and 2) improvements in technology and facility design can eliminate emissions from normally operating equipment such as pneumatic controllers (Nisbet et al., 2020). Knowing which pieces of equipment are emitting and their emission rates over time is necessary to both prioritize mitigation efforts and verify that long term emission reduction strategies are working. However, estimating emission source and rate based on ambient methane measurements is a challenging inverse problem due to the turbulent nature of near-field atmospheric transport, the degree of variability in emission rates, and the fact that only a subset of oil and gas equipment are emitting at a given time.
In this article, we analyze in situ sensor data from an experimental oil and gas site where methane is released at known rates in a way that emulates real-world emission characteristics. This site is called the Methane Emissions Technology Evaluation Center (METEC). Sensor data collected at METEC come as close as possible to data from a real oil and gas site while still knowing ground truth emission source and rate. We develop a Bayesian hierarchical model to estimate emission source and rate at METEC, and by extension, on real oil and gas sites. The model accommodates multisource scenarios and assumes that potential source locations are known a priori. We call this model the multisource detection, localization, and quantification (MDLQ) model. This is the first study, to our knowledge, that pairs a novel method for source apportionment with a controlled release experiment that is sufficiently realistic to generalize to operating oil and gas sites.
@@ 1.1 Methane releases at METEC
The METEC facility (shown in Figure 1) is instrumented with real oil and gas equipment, such as wellheads and separators, but does not actually produce oil or gas. Instead, known quantities of methane are released from typical leak points on the equipment. From February 5 to April 30, 2024, 337 controlled releases were conducted at METEC in a way that emulates the following characteristics of real oil and gas methane emissions: emissions can last from minutes to days with rates that vary over time (Daniels et al., 2024b), any subset of the equipment can emit simultaneously (Wang et al., 2022), and there can be periods of no emissions (Cusworth et al., 2021).
@@
This experiment provides a much larger and more realistic sample of methane emissions than other site-level inversion studies in the literature. For example, Cartwright et al. (2019) analyze two releases, Newman et al. (2024) analyze three releases, and Kumar et al. (2022) analyze 26 releases. The ability to apportion methane emissions using a small number of sensors is influenced by site layout, wind, and other meteorological conditions, making it important to assess the performance of a statistical inversion over a wide range of conditions.
Furthermore, the releases in this study better emulate the emission characteristics of active oil and gas sites than similar studies in the literature. For example, Weidmann et al. (2022) evaluate their method on methane releases with known start and end times, and Hirst et al. (2020) perform releases in flat, empty locations with simple atmospheric transport. As such, the modeling results in this study are more likely to generalize to real oil and gas sites of similar size and complexity to METEC and provide a baseline level of apportionment accuracy on these sites.
@@ 1.2 The continuous monitoring inverse problem
The MDLQ model uses methane concentration measurements from a network of in situ sensors placed directly on oil and gas sites to apportion emissions. We refer to this type of sensor network as a continuous monitoring system (CMS). The locations of the CMS sensors in this study are shown as pins in Figure 1. Each sensor measures ambient methane concentrations once per second, which are averaged in situ to the minute-frequency before being stored and used for all subsequent analysis. This averaging step reduces data volume while still capturing short-lived enhancements at the sub-minute scale.
Let $\boldsymbol{y}_{k}\in\mathbb{R}^{l}$ denote the minute-frequency concentration data from sensor $k$, where $l$ is the number of minutes within the temporal domain of the inversion (discussed in Section 3.3). The full observation vector is then defined as $\boldsymbol{y}=(\boldsymbol{y}_{1}^{T},...,\boldsymbol{y}_{m}^{T})^{T}\in\mathbb{R}^{n}$, where $m$ is the number of CMS sensors installed on the site and $n=l\times m$ is the total number of observations across sensors. As with most remote sensing technologies, an atmospheric inversion is required to translate $\boldsymbol{y}$ into estimates of emission source and rate.
@@
where emission rates for the $p$ potential sources on the site, $\boldsymbol{\beta}=(\beta_{1},...,\beta_{p})^{T}\in\mathbb{R}^{p}$, are assumed to be constant within the temporal domain of the inversion. The design matrix is often constructed as $X=[\boldsymbol{x}^{(1)}\cdot\cdot\cdot\boldsymbol{x}^{(p)}]\in\mathbb{R}^{n\times p}$, where each $\boldsymbol{x}^{(i)}\in\mathbb{R}^{n}$ for $i=\{1,...,p\}$ contains simulated methane concentrations from the transport model at all sensor locations assuming that source $i$ is emitting at a unitary rate. Similar to $\boldsymbol{y}$, each $\boldsymbol{x}^{(i)}$ concatenates the $l$ simulated concentrations at each of the $m$ sensors into one vector. In many studies, the errors $\boldsymbol{\epsilon}=(\epsilon_{1},...,\epsilon_{n})^{T}\in\mathbb{R}^{n}$ are assumed normal, independent, and identically distributed (Ars et al., 2017; Cartwright et al., 2019; Hirst et al., 2020; Newman et al., 2024), which is another limitation given that $\boldsymbol{y}$ are time series data. In this study, we are primarily concerned with estimates for $\boldsymbol{\beta}$, as these tell us the subset of sources that are emitting and their rate.
Our approach, implemented in the MDLQ model, assumes the same linear model as Equation 1 at the data level, but further models $\boldsymbol{\beta}$ as a mixture distribution with a point mass at zero. This prior model allows for identically zero emission rate estimates, a necessary property given that potential sources are often not emitting, and extends the method in Weidmann et al. (2022), who used a narrow Gaussian to approximate the point mass. This extension eliminates the need to tune the relative widths of the components of the mixture distribution. Finally, following Ganesan et al. (2014), we allow $\boldsymbol{\epsilon}$ to be autocorrelated since the elements of $\boldsymbol{y}$ are observed over time at 1-minute intervals.
@@ 2.1 METEC controlled release and measurement data
Methane concentration data for the METEC experiment come from a network of 10 CMS sensors installed around the facility at a height of 2 m. Eight of the sensors are also equipped with an anemometer to measure wind speed and direction in the horizontal plane. Wind data are necessary to simulate the transport of methane from the sources to the sensors (discussed further in Section 2.2). The properties (e.g., accuracy and precision) of the CMS sensors in this study are comparable to the sensors being used operationally to mitigate emissions. Full sensor specifications are provided in Section S1 of the Supporting Information (SI) file.
Unlike many studies in the literature, the controlled releases in this experiment were performed around-the-clock to evaluate the model under both daytime and nighttime meteorology. Each release consists of emissions from a subset of the five potential emission sources shown in Figure 1. Emission rates were constant within a given release, but could be different between the emitting equipment. Across the 337 releases, equipment-level release rates ranged from 0.08 to 7.2 kg/hr, emission durations ranged from 1.1 minutes to 8.0 hours, and source heights ranged from 0.91 m (West Wellheads) to 5.0 m (Tanks).
For this study, the exact source locations and heights were provided by the METEC operations team. In practice, these quantities must be estimated before performing source apportionment with the MDLQ model. Publicly available satellite imagery can be used for an accurate estimate of location (e.g., latitude and longitude) and a coarse estimate of height. More precise estimates of source height can be obtained directly from the oil and gas operator or through detailed drone imagery. For pipelines, source locations can be assigned at maintenance access points and control valves, where a large majority of pipeline emissions originate (Zimmerle et al., 2015).
@@
As an example of the METEC controlled release data, Figure 2 shows the true emission state during a 24-hour period and the corresponding concentration and wind data from the CMS sensors. In Figure 2(b), we plot the minute-by-minute maximum concentration observation across the 10 CMS sensors as a condensed summary of the data. Figure S4 in the SI shows the methane concentration measurements from all 10 CMS sensors separately. Notice that the concentration enhancements align with the timing of the controlled releases, but the amplitude of the enhancements do not follow the magnitude of the releases. This is because the sensors are not located at the release points, and meteorological effects such as wind speed and dispersion affect the transport of methane and the resulting concentrations at the sensor locations. Modeling this transport is critical for an accurate inversion of emission source and rate.
@@ 2.2 Atmospheric transport model
Consider a coordinate system in which emission source $i$ is located at $(0,0,H_{i})$. Puffs of methane from this source will all originate at this point. To avoid confusion with the symbols in Equation 1, we refer to locations in this coordinate system as $(x^{\prime},y^{\prime},z^{\prime})$. New puffs are created at a given frequency, $\delta t$, by default once per second, and the coordinates are rotated for each puff so that the positive $x^{\prime}$-direction is pointed in the downwind direction at the time of puff creation. In this coordinate system, the predicted methane concentration at sensor location $(x^{\prime},y^{\prime},z^{\prime})$ from a given puff, index by $j$, that originated at $(0,0,H_{i})$ and has been in existence for time $t$ is modeled as
$\begin{split}c_{j}^{(i)}(x^{\prime},y^{\prime},z^{\prime},t)=&\frac{Q_{j}}{(2\pi)^{3/2}\sigma_{y}^{2}\sigma_{z}}\exp{\left(-\frac{(x^{\prime}-u_{j}t)^{2}+y^{\prime 2}}{2\sigma_{y}^{2}}\right)}\\ &\left[\exp{\left(-\frac{(z^{\prime}-H_{i})^{2}}{2\sigma_{z}^{2}}\right)}+\exp{\left(-\frac{(z^{\prime}+H_{i})^{2}}{2\sigma_{z}^{2}}\right)}\right],\end{split}$ (2)
where $Q_{j}$ is the mass of methane contained in puff $j$, $\sigma_{y}$ and $\sigma_{z}$ are the dispersion parameters in the $y^{\prime}$ and $z^{\prime}$ directions, respectively, and $u_{j}$ is the wind speed at the time of puff creation. The value of $Q_{j}$ is a function of the specified emission rate, $q$, and the puff creation frequency, $\delta t$. The GPM requires time, $t$, to be discretized at a fixed multiple of the puff creation frequency, $\delta t$.
At each subsequent time step, a new puff is created at $(0,0,H_{i})$ and puffs already in existence are transported based on the wind speed and direction from their respective creation times. The GPM used in this study assumes zero advection in the vertical direction, as wind data in this dimension are hard to obtain in practice, but the puffs do grow more diffuse in all three physical dimension as they move downwind. The total methane concentration at sensor location $(x^{\prime},y^{\prime},z^{\prime})$ and time $t$ from source $i$ is then
$c^{(i)}(x^{\prime},y^{\prime},z^{\prime},t)=\sum_{j=1}^{P}c_{j}^{(i)}(x^{\prime},y^{\prime},z^{\prime},t),$ (3)
where $P$ is the number of puffs in existence at time $t$. After running the GPM, we average $c^{(i)}$ to the minute-frequency to align with the sensor data. The simulations are conducted at a higher frequency to capture the second-frequency transport dynamics that are also captured in the CMS observations.
@@ 2.3 Background estimation and removal
[…]
@@ 3.1 Bayesian hierarchical model
The MDLQ model has a hierarchical structure, summarized in Figure 4. The highest level contains a model for the observed methane concentrations. The next level contains a sparsity-inducing prior for the unobserved methane emission rates. The lowest level contains priors for all other model parameters that are not of direct interest for methane emission source apportionment but still contain important uncertainty information.
We model the CMS concentration observations as a linear combination of concentration predictions from the GPM,
$\boldsymbol{y}=X\boldsymbol{\beta}+\boldsymbol{\epsilon},$ (4)
where emission rates for the $p$ potential sources on the site, $\boldsymbol{\beta}=(\beta_{1},...,\beta_{p})^{T}\in\mathbb{R}^{p}$, are assumed to be constant within the temporal domain of the inversion, $\mathcal{D}=\{1,...,l\}$. To account for the fact that oil and gas methane emissions vary over time, we apply the MDLQ model to sequential, non-overlapping domains of length $l=30$ minutes as new CMS data become available. This windowed approach is described fully in Section 3.3; the remainder of this section defines the MDLQ model for one such temporal domain.
As in Section 1.2, we denote the background-removed, minute-frequency concentration observations from sensor $k$ as $\boldsymbol{y}_{k}\in\mathbb{R}^{l}$. We then define the full observation vector as $\boldsymbol{y}=(\boldsymbol{y}_{1}^{T},...,\boldsymbol{y}_{m}^{T})^{T}\in\mathbb{R}^{n}$, where $m$ is the number of CMS sensors installed on the site and $n=l\times m$ is the total number of observations across sensors. This formulation allows us to leverage information from all sensors jointly instead of performing a separate inversion for each.
Using the notation from Section 2.2, we denote GPM predictions at sensor $k$ as
$\boldsymbol{x}^{(i)}_{k}=(c^{(i)}(x_{k}^{\prime},y_{k}^{\prime},z_{k}^{\prime},1),...,c^{(i)}(x_{k}^{\prime},y_{k}^{\prime},z_{k}^{\prime},l))^{T}\in\mathbb{R}^{l},$ (5)
assuming that source $i$ is emitting at a unitary rate ($q$ = 1 g/s). Note that $\boldsymbol{x}^{(i)}_{k}$, and hence $\boldsymbol{\beta}$, are a linear function of $q$. However, we scale the posterior estimate of $\boldsymbol{\beta}$ by $q$, so this GPM parameter does not impact the source apportionment results. Similar to the definition of $\boldsymbol{y}$, we concatenate GPM predictions at all sensor locations into a univariate vector given by
$\boldsymbol{x}^{(i)}=({\boldsymbol{x}^{(i)}_{1}}^{T},...,{\boldsymbol{x}^{(i)}_{m}}^{T})^{T}\in\mathbb{R}^{n}.$ (6)
Finally, we define the design matrix as
$X=[\boldsymbol{x}^{(1)}\cdot\cdot\cdot\boldsymbol{x}^{(p)}]\in\mathbb{R}^{n\times p}.$ (7)
@@
We assume that the errors $\boldsymbol{\epsilon}=(\epsilon_{1},...,\epsilon_{n})^{T}\in\mathbb{R}^{n}$ follow a multivariate normal, such that $\boldsymbol{\epsilon}\sim N(0,\sigma^{2}R)$, where $R\in\mathbb{R}^{n\times n}$ is an AR(1) correlation matrix with entries $R]_{i,j}=r^{|i-j|}$ and $r\in[0,1)$ is the autoregressive coefficient or correlation parameter. Under this parameterization, $\sigma^{2}$ represents the marginal error variance rather than the innovation variance of the AR(1) process. Modeling autocorrelation in $\boldsymbol{\epsilon}$ extends previous work and results in better coverage of the credible intervals for $\boldsymbol{\beta}$, shown via a simulation study in Section S8 of the SI.
The order of the autoregressive error model was selected via a visual analysis of partial autocorrelation functions (PACFs), shown in Section S11 of the Supporting Information. Specifically, we ran the MDLQ model on the METEC data using an AR(1) error model, computed residuals $\hat{\boldsymbol{\epsilon}}=\boldsymbol{y}-X\hat{\boldsymbol{\beta}}$, and formed the implied one-step-ahead innovations $\hat{\boldsymbol{\eta}}_{t}=\hat{\boldsymbol{\epsilon}}_{t}-\hat{r}\hat{\boldsymbol{\epsilon}}_{t-1}$. We then examined PACFs of $\hat{\boldsymbol{\eta}}_{t}$ within each inversion window, noting that the innovations should be uncorrelated under a correctly specified AR(1) model. While modest residual autocorrelation remains, we do not believe that it warrants a higher autoregressive order, especially since the additional complexity could lead to identifiability issues with $\boldsymbol{\beta}$.
@@
The emission rate parameter for each source, $\boldsymbol{\beta}_{i}$ takes the form of a mixture model with a point mass at 0 kg/hr, given by
$\beta_{i}$ $\sim\begin{cases}0&z_{i}=0\\ \text{Exp}(\tau_{i}^{2})&z_{i}=1,\end{cases}$ (8)
@@
where the indicator $z_{i}\in\{0,1\}$ is modeled as $z_{i}\sim\text{Bernoulli}(\theta_{i})$. This model is an example of a sparsity enforcing spike-and-slab prior (Mitchell and Beauchamp, 1988). Sparsity in the emission rates is an important property of the MDLQ model, as emission sources on oil and gas sites are often not emitting. Previous work has found that emission rates over time follow right skewed distributions (Daniels et al., 2023; Khaliukova et al., 2025), so we use an exponential for the non-zero component of the mixture parameterized by scale $\tau_{i}^{2}$. The exponential also forces $\boldsymbol{\beta}$ to have non-negative support, which is a necessary constraint given that there are no methane sinks on oil and gas sites.
@@
This variant of the spike-and-slab prior uses a discrete point-mass as the spike rather than a continuous distribution, such as an exponential with scale parameter much smaller than $\tau_{i}^{2}$. This allows for $\beta_{i}$ to be sampled as identically zero, rather than a very small value. Furthermore, it eliminates the need to prescribe or infer the scale parameter of the narrow distribution approximating the discrete point-mass, which can be challenging to do in such a way that the MCMC sampler is efficient and mixes well.